On January 2, 2016, Alpha Corporation issued 15,000 shares of $10 par value common stock for $15 per share. On March 1, 2016, Alpha reacquired 1,000 of these shares when they were trading $20 each. Use this information to prepare the General Journal entries (without explanation) for January 2 and March 1. Format BIU General Journal: Date Accounts Debit Credit
Answer #1
Debit Credit
January 2,2016 Cash 225000 =15000*15
        Common Stock 150000 =15000*10
        Paid-in Capital in Excess of Par- Common Stock 75000
March 1,2016 Treasury Stock 20000 =1000*20
        Cash 20000